13. Multiply and simplify: (7+9i)(3+2i).
Write the answer in the form a+bi, where a and b are real numbers.

Answer by Alan3354(69443)

Multiply and simplify: (7+9i)(3+2i).
Write the answer in the form a+bi, where a and b are real numbers.
-------------------
I'll do a similar problem:
(4+5i)*(2+7i)
Multiply like any 2 binomials.
= 8 + 28i + 10i + 35i^2
= 8 + 38i + 35i^2
i^2 = -1
= 8 + 38i - 35
= -27 + 38i
**************************
You do this one.
I can check your work if you send it via the TY note.